OpenStudy (anonymous):

Solve the inequality 1/9 |5x-3|-3 ≥ 2 ??

OpenStudy (anonymous):

I got \[x \ge 53/5 \] or \[x \le -47/5\]

jimthompson5910 (jim_thompson5910):

1/9 |5x-3|-3 ≥ 2 1/9 |5x-3| ≥ 2+3 1/9 |5x-3| ≥ 5 |5x-3| ≥ 5*9 |5x-3| ≥ 45 5x-3 ≥ 45 or 5x-3 ≤ -45 I'll let you finish

OpenStudy (anonymous):

@jim_thompson5910 Ok! Thank you! :) I got x ≥ 48/5 or 5 ≤ -42/5 Do you have an idea of how I could graph that??

jimthompson5910 (jim_thompson5910):

did you mean x ≥ 48/5 or x ≤ -42/5 ?

jimthompson5910 (jim_thompson5910):

To graph this, convert each fraction to a decimal 48/5 = 9.6 -42/5 = -8.4 So x ≥ 48/5 or x ≤ -42/5 is the same as x ≥ 9.6 or x ≤ -8.4 Now draw out a number line. Plot closed circles at x = 9.6 and x = -8.4 Then shade to the left of -8.4 and to the right of 9.6 and it will look something like this